OpenStudy (anonymous):

Find three consecutive integers such that the sum of twice the smallest and three times the largest is 126.

OpenStudy (anonymous):

24 25 26

OpenStudy (anonymous):

you have been given that 2a+3(a+2)=126 so, simply solve for a: - expand the brackets, 2a+3a+6=126 collect like terms; 5a=120 divide by 5; a=24 this gives the first number in the sequence so then a+1=25 and a+2=26

OpenStudy (anonymous):

Thanks!